# Mixture of Experts Using Tensor Products

### v1

@@ Abstract
In multi-task learning, the conventional approach involves training a model on multiple tasks simultaneously. However, the training signals from different tasks can interfere with one another, potentially leading to negative transfer. To mitigate this, we propose a novel modular language model (TensorPoly), equipped with an inventory of latent modules and featuring a dynamic routing function that allocates modules to each task. Specifically, we reparameterize Low-Rank Adaptation (LoRA) by employing an entangled tensor through the use of tensor product operations, and call the resulting approach TLoRA. We tailor two innovative routing functions according to the granularity of the routing function: TensorPoly-I which directs to each rank within the entangled tensor while TensorPoly-II offers a finer-grained routing approach targeting each order of the entangled tensor. The experimental results from the multi-task T0-benchmark demonstrate that our modular LMs TensorPoly surpass the strong dense approaches, highlighting the potential of modular language models to mitigate negative inference in multi-task learning and deliver superior outcomes.
@@ 1 Introduction
Recently, the de facto paradigm for natural language understanding (NLU) tasks has centered on leveraging large language models (he2021towards) are pre-trained on a vast corpus of unlabelled data and subsequently fine-tuned for specific tasks (qiu2020pre; ye2021crossfit). While this approach has significantly advanced the field, it often requires substantial computational resources and may not efficiently transfer knowledge across diverse tasks. In addition, fine-tuning tasks independently can lead to negative transfer, where the lack of shared information across tasks, fails to achieve compositional generalization (poly).
@@
To address the aforementioned issues, there have, roughly, been two general lines of research. The first line is to mitigate the computation and memory issue, several lightweight alternatives known as parameter-efficient finetuning (PEFT) have been proposed to update only a small number of extra parameters while keeping most pre-trained parameters frozen (houlsby2019parameter; li2021prefix; lora). However, these solutions need to train each adapter for each task, which does not take into account that test tasks may require solving different combinations of sub-problems compared to training tasks (vu2020exploring), thus failing to achieve compositional generalization (rosenbaum2019routing; ponti2021inductive). Moreover, specializing the model toward different tasks during training can yield gradients that are poorly aligned, resulting in negative transfer (wang2020gradient; lialin2023scaling).
@@
The second line is to facilitate information sharing across multiple tasks using multi-task learning (MTL) approaches (caruana1997multitask; zhang2021survey; liu2019multi), which simultaneously trains the model on several tasks, allowing it to learn shared representations that benefit all tasks involved. However, MTL necessitates access to all training tasks during the training phase, meaning that incorporating new tasks requires retraining the model from scratch. This requirement significantly increases the computational burden and limits the flexibility of the model to adapt to new tasks efficiently.
@@
A promising approach to address the above issues is the adoption of modular deep learning (pfeiffer2023modular), where modules are typically implemented as parameter-efficient approaches (lora; adapter_layer; bach2022promptsource). Information flow is conditionally routed to a subset of these modules, where it is then aggregated. This design facilitates the positive transfer and systematic generalization (pfeiffer2023modular). Recently, Poly and MHR were designed to tackle diverse tasks by leveraging different combinations of latent experts (poly; mhr). Given $|\mathcal{T}|$ tasks, there are only $|\mathcal{K}|<|\mathcal{T}|$ experts are trained, we treat these $|\mathcal{K}|$ experts as latent experts, and each task-specific adapter can be obtained by a linear combination of latent experts. These approaches allow for the recombination of previously learned experts to solve test tasks more effectively, showcasing a significant advancement in task adaptability and learning efficiency. During both the multi-task pre-trained and fine-tuning, Poly/MHR concurrently optimizes the parameters of the expert inventory and a task-experts routing function, where each task-specific expert expressed by a linear combination of the activated experts. To this end, a natural question is: can we craft a non-linear function to merge experts?
Figure 1: LoRA reparamarized by an entangled tensor with a rank = 2 and order = 3. Different granularity of modules induces two routing approaches. TensorPoly-I will select the rank of the tensor while TensorPoly-II select the finer-grained order of the entangled tensor.
@@
To answer this question, we propose a variant Poly model: TensorPoly, which mixture experts using tensor product (smolensky1990tensor). The tensor product $V\times W$ of two vector spaces $\mathcal{V}$ and $\mathcal{W}$ is a vector space associated with a bilinear map $V\times W\rightarrow V\times W$3.2). This process enables the capturing of higher-order interactions and structural relationships between the input spaces (panahi2019word2ket; kye2023compositions; gan2022morphte). As depicted in Figure 1, we have reparameterized LoRA (lora) adapters by employing an "entangled" tensor structure. Consequently, the traditional training matrix $M\in\mathbb{R}^{d\times r}$ in LoRA is reparameterized into a more finer-grained tensor $\mathcal{L}\in\mathbb{R}^{N\times r\times\lceil\sqrt[N]{d}\rceil\times R)}$, named as TLoRA (4.1). This reparameterization allows for a more nuanced manipulation of the model’s parameters, facilitating a more precise and efficient adapter process. The entangled tensor configuration introduces two critical hyper-parameters: the tensor rank (R) and the tensor order (N). Leveraging these parameters, we have developed two distinct routing functions designed to select modules with varying levels of granularity. As depicted in Figure 1, TensorPoly-I employs a routing mechanism that assigns distribution scores to different tensor ranks, facilitating the selection of modules based on their rank granularity. Further advancing this concept, we propose a more refined routing function, TensorPoly-II, which targets even finer-grained tensors as activated modules. Each module is associated with a specific order of the entangled tensor. Once modules are selected via the routing function, they are aggregated through a tensor product operation, enabling a sophisticated and dynamic assembly of modular skills.
We evaluate our methods against a series of competitive baselines for few-shot task adaptations benchmark (t0_bench). Our experiments reveal several key insights:
  1. 1.
    Modular language models such as Poly and TensorPoly frameworks, consistently outperform traditional PEFT approaches LoRA and TLoRA, underscoring the effectiveness of our routing function in facilitating positive transfer across multi-task environments.
  2. 2.
    TensorPoly-I demonstrates competitive results as Poly, while simultaneously reducing the number of training and adaptation parameters required. This efficiency gain highlights the benefits of our tensorized module approach in achieving high performance with lower parameter overhead.
  3. 3.
    A comparative analysis between TensorPoly-I and TensorPoly-II indicates that the latter’s finer-grained routing mechanism does not contribute to improved performance in tensor product routing scenarios. This outcome suggests that while granularity in module selection is valuable (mhr), there is a complexity threshold beyond which additional granularity may not yield further benefits.
@@
  • We propose a novel PEFT approach TLoRA that achieves competitive results while only use $\mathcal{O}(N\times r\times\lceil\sqrt[N]{d}\rceil\times R)$ training parameters compared to LoRA $\mathcal{O}(d\times r)$, highlighting our approach’s high parameter efficiency.
  • We propose a novel modular language model TensorPoly using tensorized modules TLoRA, designed to enhance positive transfer and generalization across multi-task scenarios.
  • The findings from our T0 benchmark demonstrate that our modular language model significantly enhances the performance of a single module, underscoring the critical role of routing in scenarios involving multiple tasks.
@@ 2 Related Work
PEFT Parameter efficient fine-tuning (PEFT) methods facilitate efficient adaptation of LLMs without updating all the training parameters, thereby reducing the memory and computation cost (zhang2023llama). One kind of PEFT approach focuses on adding modules to LLMs, and only these small modules will be trained while the backbone model is kept frozen and shared across tasks. For example, Adapter Tuning inserts small neural modules (adapters) between the layers of the basic model (adapter_layer), whereas Prefix Tuning and Prompt Tuning add tunable vectors to the input or hidden layer of the base model (li2021prefix; lester2021power). Another kind of research model is the incremental update of the pre-trained weights in a parameter-efficient way, without modifying the model architecture. Bit kit fixes all training parameters and only finetunes the additive bias term (zaken2021bitfit). Diff Pruning learns a task-specific “diff vector” that extends the original pre-trained parameters. As the number of tasks increases, Diff Pruning only requires storing only a small diff vector for each task (guo2020parameter). The seminal paper lora proposes a method named LoRA that parameterizes incremental weights $\Delta$ as a low-rank matrix by the product of the down projector matrix and up projector matrix. LoRA achieves comparable or even better performance than full fine-tuning (lora). zhang2023adaptive demonstrates that weight matrices in the top layers are more important than those in the bottom layers. They propose a new method: Adaptive Low Rank Adaptation (AdaLoRA) that dynamically allocates the parameter budget among weight matrices during LoRA-like finetuning (zhang2023adaptive). AdaLoRA adjusts the rank of incremental matrices for different layers. liu2024dora introduces a new approach DoRA to investigate the inherent differences between full finetuning and LoRA. In our approach, TLoRA reparameterize LoRA using tensor products, utilizing the finer-grained tensors as modules.
@@
MTL To share the information across multiple tasks, AdapterSoup chronopoulou2023adaptersoup trains each adapter for each domain, then it performs weight-space averaging of adapters trained on different domains. (huang2023lorahub) introduce LoRAhub to aggregate the LoRA modules trained on diverse tasks. They first train a group of LoRA modules which are specialized in each task, then randomly select a subset of modules, and finally learn a set of weights to combine these LoRA models using gradient-free optimization. AdapterFusion (pfeiffer2020adapterfusion) proposes a two-stage algorithm that leverages knowledge from multiple tasks. Similarly as in LoRAhub, a group of task-specific adapters learn to encapsulate the task-specific information, and in a second stage, a fusion layer combine the trained adapters. poly introduces a variable-size module routing mechanism, Poly, predicated on the notion that each task correlates with a specific subset of latent skills drawn from a comprehensive inventory of modules. Concurrently, both the modules and the routing function are trained, facilitating a more dynamic and task-specific approach to model learning. Building upon Poly, (mhr) introduces a finer-grained multi-head routing function MHR where the experimental findings underscore the significance of the routing function during the pre-training phase. Remarkably, by solely fine-tuning the routing function, MHR attains competitive results, demonstrating exceptional parameter efficiency.
Model Merging Model merging can be seen as a parameter level aggregation, combining multiple models into one single model in parameter space without access to data (matena2022merging). The simplest operation of merging is averaging the weights of different models. wortsman2022model shows that averaging the weights of multiple models finetuned with different hyperparameter configurations often improves accuracy and robustness. matena2022merging takes the aggregation function maximizing the joint likelihood of the posteriors of the model’s parameters and proposes a “Fisher merging” technique. yadav2023ties proposes a TIES-Merging method to address interference due to redundant parameter values and disagreement on the sign of a given parameter’s values across models. ilharco2022editing edits models with task arithmetic operations. jin2022dataless propose a dataless knowledge fusion method (Regression Mean) that merges models in their parameter space.
@@ 3 Background
In multiple transfer learning tasks , we define a set of tasks as $\mathcal{T}=\{\mathcal{T}_{1},...,\mathcal{T}_{|\mathcal{T}|}\}$. This set is divided into two subsets train $\mathcal{T}_{train}$ and test $\mathcal{T}_{test}$. The goal of multi-task transfer learning is to apply the knowledge from the training tasks $\mathcal{T}_{train}$ to the test tasks within $\mathcal{T}_{test}$. This process involves two main phases. Building on top of a foundation model, the first phase consists of multi-task pre-training using the dataset in tasks $\mathcal{T}_{train}$. The second consists of a few-shot adaptation, where the learned adapters are fine-tuned independently on each test task in $\mathcal{T}_{test}$. We follow the procedure from (t5) and formulate each task as a text-to-text problem, enabling standard maximum-likelihood training with teacher forcing (bengio2015scheduled) and a cross-entropy loss.
@@ 3.1 Module: LoRA
LoRA are recently proposed adapter architecture that achieve a competitive balance between performance and parameter efficiency (lora; mahabadi2021parameter). For each linear transformation corresponding to the query ($q$), key ($k$), value ($v$) and output ($o$) of the self-attention layers, LoRA modifies the base model parameters as follows:
@@
where $W_{0}$ are the (frozen) weights of the pre-trained model (e.g. T5 (t5)). $A,B\in\mathbb{R}^{d\times r}$ are low-rank learnable parameters and $s\geq 1$ is a tunable scalar hyperparameter. We schematize LoRA in Figure 2.

### v2

# Mixture of Latent Experts Using Tensor Products

@@ Abstract
In multi-task learning, the conventional approach involves training a model on multiple tasks simultaneously. However, the training signals from different tasks can interfere with one another, potentially leading to negative transfer. To mitigate this, we propose a novel latent-expert approach (TensorPoly), that balances parameter efficiency with nuanced routing methods. For experts, we reparameterize Low-Rank Adaptation (LoRA) by employing an entangled tensor through the use of tensor product operations and name the resulting approach TLoRA. For routing function, we tailor two innovative routing functions according to the granularity: TensorPoly-I which directs to each rank within the entangled tensor while TensorPoly-II offers a finer-grained routing approach targeting each order of the entangled tensor. The experimental results from the multi-task T0-benchmark demonstrate that: 1) all latent-expert approaches surpass the corresponding dense approaches, highlighting the potential of modular language models to mitigate negative inference in multi-task learning and deliver superior outcomes. 2) TensorPoly-I achieves higher parameter efficiency in adaptation and outperforms other modular LMs, which shows the potential of our approach in multi-task transfer learning  [^1] The code is released: https://github.com/microsoft/mttl.
@@ 1 Introduction
Recently, the de facto paradigm for natural language understanding (NLU) tasks has centered on leveraging large language models (LLMs) (He et al., 2021) that are pre-trained on a vast corpus of unlabelled data and subsequently fine-tuned for specific tasks (Qiu et al., 2020; Ye et al., 2021). While this approach has significantly advanced the field, it often requires substantial computational resources and may not efficiently transfer knowledge across diverse tasks. In addition, fine-tuning tasks independently would lead to negative transfer, where the lack of shared information across tasks makes it difficult to achieve compositional generalization (Ponti et al., 2023).
Figure 1: Left: Comparison between the dense models (LoRA, TLoRA) and latent-expert approaches (Poly, MHR, TensorPoly-I, TensorPoly-II). Poly/MHR use LoRA as the modules, TensorPoly-I and TensorPoly-II use TLoRA as the modules. Right: Adaptation parameters across different approaches in the fine-tuning process.
@@
To address the aforementioned issues, there have been, roughly, two general lines of research. The first line aims to mitigate the computation and memory issue using lightweight alternatives known as parameter-efficient fine-tuning (PEFT), which updates only a small number of extra parameters while keeping most pre-trained parameters frozen (Houlsby et al., 2019b; Li & Liang, 2021; Hu et al., 2021). However, these solutions need to train an adapter for each task, which does not take into account the fact that test tasks may require solving different combinations of sub-problems compared to training tasks (Vu et al., 2020), thus failing to achieve compositional generalization (Rosenbaum et al., 2019; Ponti, 2021).
@@
The second line is to facilitate information sharing across multiple tasks using multi-task learning (MTL) approaches (Caruana, 1997; Zhang & Yang, 2021; Liu et al., 2019), which simultaneously train the model on several tasks, allowing it to learn shared representations for all tasks involved. However, MTL necessitates access to all training tasks during the training phase, meaning that incorporating new tasks requires retraining the model from scratch. This requirement significantly increases the computational burden and limits the flexibility of the model to adapt to new tasks efficiently.
@@
A promising approach to address the above issues is the adoption of modular language models (Pfeiffer et al., 2023; Ponti et al., 2023; Caccia et al., 2023), where modules are typically implemented as PEFTs for different tasks (Hu et al., 2021; Houlsby et al., 2019a; Bach et al., 2022). Information flow is conditionally routed to a subset of these modules, which are then aggregated for the given task. This design facilitates the positive transfer and systematic generalization (Pfeiffer et al., 2023). Recently, Poly and MHR were designed to handle diverse tasks by leveraging different combinations of latent experts (Ponti et al., 2023; Caccia et al., 2023). Given $|\mathcal{T}|$ tasks, there are only $|\mathcal{K}|<|\mathcal{T}|$ experts trained. We treat these $|\mathcal{K}|$ experts as latent experts, and each task-specific adapter can be obtained by a linear combination of latent experts. During both the multi-task pre-trained and fine-tuning, Poly implements adapters with LoRAs and concurrently optimizes the LoRA inventory and a routing function. MHR partitions the LoRAs into multiple heads and uses a finer-trained routing among these heads. However, LoRA adapters are still limited in parameter efficiency, especially with expert libraries involving a huge number of adapters. In addition, we notice that previous approaches only use linear combinations of experts, which implicitly assume that the given task has a linear relationship with the expert modules, whereas the relations could be much more complicated in practice. To this end, the question of developing a modular language model that balances parameter efficiency with complex routing methods is critical for advancing the scalability and functionality of multi-task transfer learning.
@@
To answer this question, we devise a new variant Poly model: TensorPoly, which mixes experts using tensor product (Smolensky, 1990), an operation that maps two vectors in spaces $\mathcal{V}$ and $\mathcal{W}$ to a vector in the tensor product space $\mathcal{V}\otimes\mathcal{W}$ of the two vector spaces, associated with a bilinear map $\mathcal{V}\times\mathcal{W}\rightarrow\mathcal{V}\otimes\mathcal{W}$3.3). This process enables the capture of higher-order interactions and structural relationships between input spaces (Kye, 2023). Tensor product has been successful as a strategy for compressing word embeddings (Panahi et al., 2019; Gan et al., 2022). To achieve a higher parameter-efficient adapter, we have reparameterized LoRA (Hu et al., 2021) adapters by employing an “entangled” tensor structure, where the high hidden and intermediate dimensions are decomposed into smaller dimensions in tensor product form. Consequently, the training matrix $M\in\mathbb{R}^{d\times r}$ in LoRA is reparameterized into a finer-grained tensor $\mathcal{L}\in\mathbb{R}^{N\times r\times\lceil\sqrt[N]{d}\rceil\times R)}$, named TLoRA (4.1). This reparameterization allows for a more nuanced manipulation of the model’s parameters, facilitating a more efficient adapter process for more complex tasks. The entangled tensor configuration introduces two critical hyper-parameters: the tensor rank (R) and the tensor order (N). Leveraging these parameters, we develop two distinct routing functions designed to select modules on varying levels of granularity. As depicted in Figure 3, TensorPoly-I employs a routing mechanism that assigns distribution scores to different tensor ranks, facilitating the selection of modules based on their rank granularity. Further advancing this concept, we propose a more refined routing function, TensorPoly-II, which targets even finer-grained tensors as activated modules. Each module is associated with a specific order of the entangled tensor. Once modules are selected via the routing function, they are aggregated through a tensor product operation, enabling a sophisticated and dynamic assembly of modular skills.
To evaluate the effectiveness and parameter efficiency of our approach, we apply our methods against a series of competitive baselines on T0 (Sanh et al., 2021), a widely used benchmark in multi-task transfer learning covering a high variety of language understanding tasks. Our experiments reveal several key insights: i) Modular language models such as Poly, MHR and TensorPoly frameworks consistently outperform traditional PEFT approaches LoRA and TLoRA (Figure 1), underscoring the effectiveness of modular LLMs in facilitating positive transfer across multi-task environments. ii) TensorPoly-I demonstrates competitive results against Poly and MHR, while simultaneously achieving higher parameter efficiency in adaptation. This efficiency gain highlights the benefits of our tensorized module approach in achieving high performance with lower parameter overhead. iii) A comparative analysis between TensorPoly-I and TensorPoly-II indicates that the latter’s finer-grained routing mechanism does not contribute to improved performance. This outcome suggests that while granularity in module selection is valuable (Caccia et al., 2023), there is a complexity threshold beyond which additional granularity may not yield further benefits.
@@
  • We propose TLoRA that achieves competitive results while only using $\mathcal{O}(N\times r\times\lceil\sqrt[N]{d}\rceil\times R)$ training parameters compared to LoRA $\mathcal{O}(d\times r)$, highlighting our approach’s high parameter efficiency.
  • We propose a novel modular LM TensorPoly, which balances parameter efficiency with tensor product routing. The evaluation results on T0 benchmark demonstrate that TensorPoly-I can surpass other strong modular LMs, underscoring the critical role of tensor product routing in scenarios involving multi-task transfer learning.
  • By only fine-tuning the routing function. TensorPoly-I can surpass the TLoRA with only 8.6k training parameters, achieving extremely parameter-efficient fine-tuning.
@@ 2 Related Work
Parameter-efficient fine-tuning (PEFT) methods facilitate efficient adaptation of LLMs without updating all the training parameters, thereby reducing the memory and computation cost (Zhang et al., 2023b). One kind of PEFT approach focuses on adding modules to LLMs, and only these small modules will be trained while the backbone model is kept frozen and shared across tasks. For example, Adapter Tuning inserts small neural modules (adapters) between the layers of the basic model (Houlsby et al., 2019a), whereas Prefix Tuning and Prompt Tuning add tunable vectors to the input or hidden layer of the base model (Li & Liang, 2021; Lester et al., 2021). Another kind of research model is the incremental update of the pre-trained weights in a parameter-efficient way, without modifying the model architecture. Bitfit fixes all training parameters and only fine-tunes the additive bias term (Zaken et al., 2021). Diff Pruning learns a task-specific “diff vector” that extends the original pre-trained parameters. As the number of tasks increases, Diff Pruning only requires storing a small diff vector for each task (Guo et al., 2020). The seminal paper Hu et al. (2021) proposes a method named LoRA that parameterizes incremental weights $\Delta$ as a low-rank matrix by the product of the down projector matrix and up projector matrix. LoRA achieves comparable or even superior performance to full fine-tuning (Hu et al., 2021). Zhang et al. (2023a) demonstrates that weight matrices in the top layers are more important than those in the bottom layers. They propose Adaptive Low Rank Adaptation (AdaLoRA), a new method that dynamically allocates the parameter budget among weight matrices during LoRA-like fine-tuning (Zhang et al., 2023a). AdaLoRA adjusts the rank of incremental matrices for different layers. Liu et al. (2024) introduces a new approach DoRA to investigate the inherent differences between full fine-tuning and LoRA. Baziotis et al. (2022) finds hyper-adapters are more parameter efficient than regular adapters, reaching the same performance with up to 12 times less parameters.
Compared to the works above, our approach TLoRA takes a novel perspective by reparameterzing LoRA using tensor products and utilizing finer-grained tensors as modules.
Multi-task Learning
@@
The key to MTL is information sharing across tasks. For this purpose, AdapterSoup Chronopoulou et al. (2023) trains each adapter for each domain, and performs weight-space averaging of adapters trained on different domains. Huang et al. (2023) introduce LoRAhub to aggregate the LoRA modules trained on diverse tasks. They first train a group of LoRA modules that are specialized in each task, then randomly select a subset of modules, and finally learn a set of weights to combine these LoRA models using gradient-free optimization. AdapterFusion (Pfeiffer et al., 2020) proposes a two-stage algorithm that leverages knowledge from multiple tasks. Similarly to LoRAhub, a group of task-specific adapters learn to encapsulate the task-specific information, and in the second stage, a fusion layer combines the trained adapters. Ponti et al. (2023) introduces a variable-size module routing mechanism, Poly, based on the assumption that each task correlates with a specific subset of latent skills drawn from a comprehensive inventory of modules. Building upon Poly, (Caccia et al., 2023) introduces a finer-grained multi-head routing function MHR where the experimental findings underscore the significance of the routing function during the pre-training phase. Mixture-of-expert (MoE) methods such as TaskMoE Kudugunta et al. (2021) learn a routing function that allocates modules to tasks end-to-end.
We propose TensorPoly-I and TensorPoly-II, two novel routing mechanisms based on TLoRA. TensorPoly-I is a variant of Poly using TLoRA as the modules. In this setting, each rank of the entangled tensor corresponds to a separate expert. TensorPoly-II is a finer-grained routing function targeting each order of the entangled tensor.
Expert Merging
Once the experts are activated in the forward pass, we need to aggregate their outputs. There is an increasing focus on aggregating adapters from different domains through expert merging. The simplest operation of merging is averaging the weights of different experts, where the weight of each expert is set according to the routing probability generated by the router (McMahan et al., 2017; Choshen et al., 2022; Matena & Raffel, 2022; Chronopoulou et al., 2023; Huang et al., 2023; Muqeeth et al., 2023; Ostapenko et al., 2023; 2024). Poly Ponti et al. (2023) uses the latent experts and integrates the experts by averaging the weights. MHR Caccia et al. (2023) partitioned the LoRA experts into different heads, which are eventually concatenated to obtain a merged expert.
We devise two tensor product routing functions TensorPoly-I and TensorPoly-II. Once each expert is activated, a routing function will aggregate the expert weights as an entangled tensor.
@@ 3 Background
In multiple transfer learning tasks, we define a set of tasks as $\mathcal{T}=\{\mathcal{T}_{1},...,\mathcal{T}_{|\mathcal{T}|}\}$. This set is divided into two subsets train $\mathcal{T}_{train}$ and test $\mathcal{T}_{test}$. The goal of multi-task transfer learning is to apply the knowledge from the training tasks $\mathcal{T}_{train}$ to the test tasks within $\mathcal{T}_{test}$. This process involves two main phases. Building on top of a foundation model, the first phase consists of multi-task pre-training using the dataset in tasks $\mathcal{T}_{train}$. The second consists of a few-shot adaptation, where the learned adapters are fine-tuned independently on each test task in $\mathcal{T}_{test}$. We follow the procedure from (Raffel et al., 2020) and formulate each task as a text-to-text problem.
@@ 3.1 Module: LoRA
Lora is a recently proposed adapter architecture that achieves a competitive balance between performance and parameter efficiency (Hu et al., 2021; Mahabadi et al., 2021). For each linear transformation corresponding to the query ($q$), key ($k$), value ($v$), and output ($o$) of the self-attention layers, LoRA modifies the base model parameters as follows:
@@
where $W_{0}$ are the (frozen) weights of the pre-trained model (e.g. T5 (Raffel et al., 2020)). $A,B\in\mathbb{R}^{d\times r}$ are low-rank learnable parameters and $s\geq 1$ is a tunable scalar hyperparameter. We schematize LoRA in Figure 3.
3.2 Polytropon (Poly): Mixture of Latent Experts with Linear Combination
Figure 2: Compare with three training paradigms in multi-task transfer learning. Left is the private training, for each task, we train the corresponding expert individually. Middle is the shared version, for all the tasks, we train an expert continually, and as a result, we only get one expert. Right is the latent experts model, for all the tasks, we train a subset of "latent" experts, so each corresponding expert can be seen as a linear combination of these latent experts.
Poly/ MHR addresses the multi-task problem by softly sharing latent experts across tasks. Each Poly layer contains 1) an inventory of latent experts =$\mathcal{M}=\{\phi_{1},\ldots,\phi_{m}\}$ with $|\mathcal{M}|\ll|\mathcal{T}|$; 2) a routing function $r(\cdot)$ that chooses which subset of the experts to combine for each task. Each latent expert corresponds to a LoRA adapter, where $\phi_{i}$ are its associated parameters $A^{(i)},B^{(i)}\in\mathbb{R}^{d\times r}$. $r(\cdot)$ is implemented as a task–module routing matrix $Z\in\mathbb{R}^{|\mathcal{T}|\times|\mathcal{M}|}$. $z_{\tau}=Z_{\tau,:}\in\mathbb{R}^{|\mathcal{M}|}$ is a routing vector of task ${\mathcal{T}_{\tau}}$, with cell $Z_{\tau,j}$ being the probability logits of using module $\phi_{j}$ for task $\mathcal{T}_{\tau}$ in the current layer. Differently from mixture-of-experts (Fedus et al., 2022), which perform token-level top-$k$ routing, $Z$ converges to a binary matrix, defining a soft partition over modules. This is achieved by using a Gumbel-sigmoid distribution (Jang et al., 2017) during training, with $\hat{Z}_{\tau,j}\sim\texttt{Gumbel}(Z_{\tau,j})$. At each forward pass, Poly can be defined as:
$A^{\tau}=\sum_{i}\alpha_{i}A^{(i)};\,B^{\tau}=\sum_{i}\alpha_{i}B^{(i)}$ (Poly)